OpenStudy (wondermath):

How do you prove that the slope of f(x)=c is 0 using the difference quotient? I tried it and got 1.

OpenStudy (jamesj):

The slope is \[ m = \frac{y_2 - y_1}{x_2 - x_1} = \frac{f(x_2) - f(x_1)}{x_2 - x_1} = \frac{c - c}{x_2 - x_1} = ... what? \]

OpenStudy (wondermath):

can you show me using the difference quotient ?

OpenStudy (jamesj):

This is the difference quotient: \[ \frac{f(x_2) - f(x_1)}{x_2 - x_1} \]

OpenStudy (wondermath):

really? i thought it was

OpenStudy (wondermath):